\begin{thebibliography}{99}

%\bibitem{Berrone} L.R. Berrone, C.D. Galles, Comment on “Using model systems to demonstrate instances of mathematical
%inequalities”, Am. J. Phys. 66 (87-88) (1998).
%
%\bibitem{Landsberg2} P. T. Landsberg and J. E. Pecaric, Thermodynamics, inequalities and
%negative heat capacities,  Phys. Rev. A 35, 4397–4403 (1987).
%

%
%\bibitem{Mungan2017} C. E. Mungan CE, Damped oscillations of a 
%frictionless piston in an adiabatic cylinder enclosing an ideal 
%gas. Eur. J. Phys. \textbf{38}, 035102 (2017).
%

%
%\bibitem{Mungan2015} C. E.Mungan, Entropy change for the irreversible heat transfer between two finite objects, Eur. J. Phys. \textbf{36}, 048004 (2015).
%
%\bibitem{kiatgamolchai2015} S. Kiatgamolchai, A graphical proof of the positive entropy change in heat transfer between
%two objects, Phys. Teach. \textbf{53}, 95–6 (2015).
%
%\bibitem{Lima} F.M.S. Lima, Eur. J. Phys. \textbf{36}, 068001 (2015).
%
%\bibitem{Freire} W. H. C. Freire, D.  Leite. (2016). Variação de entropia na termalização espontânea de dois blocos distintos revisitada, Revista Brasileira de Ensino de Física, \textbf{38}(1), 1702 (2016).
%		
%\bibitem{Maligranda} L. Maligranda, The AM-GM Inequality is Equivalent to the Bernoulli Inequality,
%The Mathematical Intelligencer \textbf{34} 1-2 (2012).
%
%\bibitem{Halliwell} J. J. Halliwell, J. P\'erez-Mercader and
% W. H. Zurek, eds., \emph{Physical Originsof Time Asymmetry} 
%(Cambridge Univ. Press., Cambridge, 1994).




\bibitem{Talwalkar}https://mindyourdecisions.com/blog/2013/08/05/monday-puzzle-what-is-greater-epi-or-pie/ (visited on 25/08/2023).


%\bibitem{Gislason} E. A. Gislason, N. C.  Craig, Pressure-volume integral
%expressions for work in irreversible processes., J. Chem. Educ.
%\textbf{84}, 499 (2007).
%
%\bibitem{Tait} P. G. Tait, Physical proof that the geometric mean of any number of
%quantities is less than the arithmetic mean, Proc. Roy. Soc. Edin. 6
%(1868/9) p. 309; reprinted in Scientific Papers, Vol. 1 (1898) p. 83. 
%
%\bibitem{Sommerfeld} A. Sommerfeld, Thermodynamics and statistical mechanics (Trans. J.
%Kestin), Academic Press, New York (1964), German original 1952.
%
%
%
%%\bibitem{Wang} L. Wang, Second law of
%%thermodynamics and arithmetic-mean–geometric-mean inequality, Internat. J. Modern Phys.
%%B, \textbf{13}(21-22):2791–2793 (1999).
%
%\bibitem{Graham} C. Graham, T. Tokieda, 
%An Entropy Proof of the Arithmetic Mean–Geometric Mean Inequality, 
%The American Mathematical Monthly \textbf{127}:6, 545-546 (2020).
%
%\bibitem{vanWylen} G. J. Van Wylen G.J. and R.E Sonntag, 
%Fundamentals of Classical Thermodynamics, 1st edition, 
%John Wiley and Sons, New York, 1965. 
%
%\bibitem{Cengel} Y.A. Cengel, and M.A. Boles, 
%\emph{Thermodynamics: An Engineering Approach}, 4th edition,
%1056 pp. McGraw-Hill, New York (2001).
%
%\bibitem{Callen} H. B. Callen, \emph{Thermodynamics and an Introduction to Thermostatistics}, 2nd edition
%(John Wiley, 1985).
%
%\bibitem{Moran} M. J. Moran, H.N. Shapiro, \emph{Fundamentals of engineering thermodynamics} (John Wiley, New York, 1988).
%
%\bibitem{Thomas}C. Thomas, M. Badger, E. Ventura-Medina and C. Sangwin, 
%Puzzle-based learning of mathematics in engineering, 
%Engineering Education \textbf{8}(1), 122-134 (2013). 
 
\bibitem{Pickover}C. A. Pickover, Pickover, 
Wonders of Numbers (Oxford: Oxford University Press, 2001).

\bibitem{Note1} Note that, since the process is out of equilibrium, the pressure 
exerted on the gas is different from the gas pressure, which is probably 
not even well-defined during the whole process.

\bibitem{Mungan} C. E. Mungan, Damped oscillations of a 
frictionless piston in an adiabatic cylinder enclosing an ideal 
gas. Eur. J. Phys. \textbf{38}, 035102 (2017).

\bibitem{Abriata} J. P. Abriata, 
Comment on a thermodynamical proof of the inequality between arithmetic and geometric mean, 
Phys. Lett. A \textbf{71}, 309–310 (1979).

\bibitem{Deakin}M. A. B. Deakin and G. J. Troup, 
The logical status of thermodynamic proofs of mathematical theorems, 
Phys. Lett. A \textbf{83} 6, 239–240 (1981).

\bibitem{Sidhu} S. S. Sidhu, 
On thermodynamic proof of mathematical results, 
Phys. Lett. A \textbf{76}, 107–108 (1980).

\bibitem{Cashwell} E. D. Cashwell and C. J. Everett, The means of order t and the laws of
thermodynamics, Am. Math. Monthly 74, 271–274 (1967).

\bibitem{Landsberg} P. T. Landsberg, A thermodynamic proof of the inequality between
arithmetic and geometric means, Phys. Lett. 67A (1978) p. 1. 

\bibitem{Tykodi} R. J. Tykodi, Using model systems to demonstrate instances of mathematical inequalities, Am. J. Phys. 64 5, 644–648 (1996).

%\bibitem{Graham} C. Graham, T. Tokieda, 
%An Entropy Proof of the Arithmetic Mean–Geometric Mean Inequality, 
%The American Mathematical Monthly \textbf{127}:6, 545-546 (2020).

\bibitem{Plastino} A. R. Plastino, A. Plastino, G. H Miller, 
Thermodynmic paths to Jensen's inequality, 
Am. J. Phys. \textbf{65}(11), 1102–1105 (1997).

\bibitem{Vallejo} A. Vallejo, 
Finding and improving bounds of real functions by thermodynamic arguments, 
Am. J. Phys. \textbf{91}(5), 366-370 (2023).


\end{thebibliography}

\RequirePackage[2020-02-02]{latexrelease}
\documentclass[twocolumn]{revtex4}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{color}
\usepackage{epstopdf}
\usepackage{float}
\usepackage{siunitx}
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}

\begin{document}
\title{Which is greater: $\boldsymbol{e^\pi}$ or $\boldsymbol{\pi^{e}}$? An unorthodox solution to a classic puzzle}
\author{Andr\'es Vallejo, Italo Bove}
\affiliation{\begin{small} Facultad de Ingenier\'{\i}a, Universidad 
de la Rep\'ublica, Montevideo, Uruguay\end{small}}
\date{\today}
\begin{abstract}
The question of the title is a famous puzzle in the field of recreational mathematics, 
and can be addressed by several approaches. A compilation of solutions, some of them very 
ingenious, can be found in \cite{Talwalkar}. In this contribution we present an alternative
solution based on the second law of thermodynamics. The method can be extended  
to derive a more general result involving the exponential function.
\end{abstract}

\maketitle

%
It is a well-known fact that when two bodies $A$ and $B$ at different temperatures 
are placed in contact, energy flows between them until thermal equilibrium is reached. 
Let us suppose that $A$ is an
incompressible solid with constant heat capacity $C$, initially at temperature $T_{1}=\pi$, and 
$B$ is a large thermal reservoir at temperature  $T_{B}=e$, where $e$ is Euler's number 
(both temperatures are measured in the same absolute temperature scale). We will 
assume that each of the systems only exchanges heat with the other.

Since there are no thermometers with infinite precision, the reader may argue that it is not 
possible to guarantee that the temperatures correspond precisely to these irrational numbers. 
However, for macroscopic systems with a large number of degrees of freedom, the temperature is 
usually considered as a continuous variable, that, a priori, can adopt any real value. In any case, 
if truncated expressions are employed, a single number after the comma is enough to obtain the 
correct relation between the exponentials in the title. 

Since the reservoir is large, the equilibrium is reached when the temperature of 
the solid equals that of the reservoir, which remains essentially constant. 
The entropy variation of the solid at the end of the process is:
	\begin{equation}\label{dS_S}
	\Delta S^{A}=C\log\left(\dfrac{T_{2}}{T_{1}}\right)=C(1-\log(\pi)).
	\end{equation}
The reservoir exchanges energy in an amount that is opposite to the 
solid's internal energy variation:
	\begin{equation}
	Q^{B}=-Q^{A}=-\Delta U^{A}=C(T_{1}-T_{2})=C(\pi-e),
	\end{equation}
so the entroy of the reservoir changes in the amount:
	\begin{equation}\label{dS_R}
	\Delta S^{B}=\dfrac{Q^{B}}{T^{B}}=C\left(\dfrac{\pi}{e}-1\right),
	\end{equation}
Finally, from Eqs. (\ref{dS_S}) and (\ref{dS_R}), we have that the global entropy change 
associated with the thermalization process is:
\begin{equation}\label{dS_U}
\Delta S^{Univ}=\Delta S^{A}+\Delta S^{B}=C\left[\dfrac{\pi}{e}-\log(\pi)\right]
\end{equation}

The key point is to note that, according to the second law of thermodynamics, the total entropy variation in any physical process must be non negative. Using that $C>0$, from Eq. (\ref{dS_U}) we obtain that
	\begin{equation}\label{ex}
	\dfrac{\pi}{e}-\log(\pi)\geq 0 \Longrightarrow \pi\geq\log(\pi^e)\Longrightarrow e^\pi\geq \pi^e,
	\end{equation}
so we conclude that $e^{\pi}$ is the greater of the two numbers. It is known as \textit{Gelfond's constant}, and it is a trascendental number whose approximate value is $e^{\pi}\simeq 23.14069$. On the other hand, $\pi^{e}\simeq 22.45915$, and it is unknown whether or not it is trascendental \cite{Pickover}.

We can formulate an alternative thermodynamic derivation of this result by considering a 
cylinder-piston device with diathermal walls containing a perfect gas in equilibrium in a
thermal bath. If the mass of the piston is such that the initial pressure in some unit 
system is $P_1=e$, and we suddenly increase the 
mass of the piston in such a way that the pressure exerted on the gas (in the same unit 
system) adopts the value $\pi$ \cite{Note1}, after transient oscillations, the system will 
reach a new equilibrium state with the bath, at pressure $P_2=\pi$ and at the initial 
temperature \cite{Mungan}. The interested reader may verify that the entropy analysis 
of this process leads to:
	\begin{equation}\label{DeltaS_univ}
	\Delta S^{Univ}=R\left[\dfrac{\pi}{e}-\log(\pi)\right],
	\end{equation}
from which the same conclusion is derived. 
It is also interesting to note that, if the final pressure (or, in the previous example, the initial temperature) is some arbitrary positive value $x$, an analogous reasoning allows to infer the following more general inequality:

	\begin{equation}\label{ineq1}
	e^x\geq x^e,\hspace{0.2cm}x\geq 0.
	\end{equation}

The possibility of deriving an inequality involving these famous mathematical 
constants by means of a physical law produces surprise and arouses interest in 
students. It usually motivates them to start exploring thermodynamic paths to generate 
other algebraic inequalities, which results in an improvement in students' skills in 
performing entropy analysis. 

Another interesting aspect that can be discussed in classroom is 
the logical status of this type of derivation. Since the results obtained are 
mathematical truths, and, therefore, independent of any physical considerations, 
it is interesting to discuss whether the thermodynamic derivation is, in fact, a proof  
in the mathematical sense. We refer the readers to Refs. \cite{Abriata,Deakin,Sidhu} 
to delve deeper into this issue. 

Instructors interested in this approach can find thermodynamic derivations 
of other famous inequalities in Refs. \cite{Cashwell,Landsberg,Tykodi} 
(inequalities between means), \cite{Plastino} (Jensen's inequality), or \cite{Vallejo} 
(Bernoulli's inequalities and bounds for the logarithmic function).


\section*{Acknowledgments}
This work was partially supported by Agencia Nacional de Investigaci\'on e Innovaci\'on and Programa de Desarrollo de las Ciencias B\'asicas (Uruguay). 


\end{document}